an angle's measure is 5 degrees less than 3 times the measure of its supplement. Find the measure of the angle and its supplement?

Answer by Leaf W.(135)

*x = supplement of "an angle"
*3x = 3 times the measure of the supplement
*3x - 5 = 5 degrees less than 3 times the measure of its supplement = "an angle" itself
*since supplementary angles are angles that add up to 180 degrees, the sum of the measures of "an angle" and its supplement must be 180; with x as the supplement and 3x - 5 as "an angle," the equation would be (3x - 5) + x = 180, which is the same as 3x - 5 + x = 180
*add up the like terms (x's): 4x - 5 = 180
*add 5 to both sides: 4x = 185
*divide both sides by 4: x = 46.25
*since x is the supplement, THE SUPPLEMENT IS EQUAL TO 46.25 DEGREES
*plug this value for x into the expression for the angle, 3x - 5: 3(46.25) - 5
*multiply 3 and 46.25: 138.75 - 5
*subtract 5 from 138.75: 133.75
*therefore, THE ANGLE IS EQUAL TO 133.75 DEGREES
Hope this helps! Good luck! =)